OpenStudy (anonymous):

Peacewise Function: f(x)=12 if x<-3 15 if -3<=x<6 20 if x>= 6

OpenStudy (anonymous):

Can someone explain it step by step please? :)

OpenStudy (jdoe0001):

hmm explain? what do you need from \(\large f(x) \begin{cases} 12\qquad &x<3 \\ \quad \\ 15\qquad &-3\le x<6 \\ \quad \\ 20\qquad & x\ge 6 \end{cases}\quad ?\)

OpenStudy (anonymous):

Evaluate the piecewise function for x = -1 and x = 3 it says

OpenStudy (jdoe0001):

hmmm I missed a sign there... anyhow \(\large f(x) \begin{cases} 12\qquad &x<-3 \\ \quad \\ 15\qquad &-3\le x<6 \\ \quad \\ 20\qquad & x\ge 6 \end{cases}\)

OpenStudy (anonymous):

So am I supposed to plug something in ?

OpenStudy (anonymous):



OpenStudy (anonymous):

f(x)=15 for x=-1 to 3

OpenStudy (anonymous):

f(-1)=15 f(3)=15

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

you got it.

OpenStudy (anonymous):

you take any value between 13 and 6 f(x)=15

OpenStudy (anonymous):

correction -3

OpenStudy (anonymous):

at x=-3,it is 15 at x=6 ,it is 20